2. + -12 points PSE64P. 041 My Notes A river has a steady speed of 0.500 m/s. A student swims upstream a distance of 1.65 km and swims back to the starting point. If the student can swim at a speed of 1.00 m/s in still water, how long does the trip take (swimming up and down in the current)? If the water were still, by how much would the trip be longer or shorter? (Give a negative answer for a shorter trip.)If the number is a decimal please include 4 digits past the decimal point if available.
Answer #1

answer) when the student goes upstream his relative speed is =1-0.5=0.500m/s

so t=distance/speed=1650m/0.500m/s=3300s

when going downstream the speed of the student is =1.00+0.500=1.500m/s

so t=1650/1.5=1100s

so the whole trip takes=3300+1100=4400s=answer

in still water the total time=3300/1=3300s

the difference=3300-4400= -1100s=answer